OpenStudy (anonymous):

Is the problem 24(x - 1) = -4(6 - x) + 20x a conditional equation, contradiction, or identity? I already solved it, the answer is x = 0.

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

\( 24(x - 1) = -4(6 - x) + 20x \)

ganeshie8 (ganeshie8):

distribute, wat do u get ?

OpenStudy (anonymous):

24x - 24 = -24 + 16x - 16x 8x - 24 = -24 + 24 8x = 0 / 8 x = 0

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

there is a mistake in ur first step, right side u will not get 16x

OpenStudy (anonymous):

Oh I forgot about the -x lol

OpenStudy (anonymous):

So.. 24x - 24 = -24 + 20x - 20x 4x - 24 = -24 + 24 4x = 0 / 4 x = 0 ?

ganeshie8 (ganeshie8):

\(\large 24(x - 1) = -4(6 - x) + 20x \) \(\large 24x - 24 = -24 + 4x + 20x\) \(\large 24x - 24 = -24 + 24x\)

OpenStudy (anonymous):

-24 = -24

OpenStudy (anonymous):

Is it identity?

ganeshie8 (ganeshie8):

Yup !!!!

ganeshie8 (ganeshie8):

you got -24 = -24 which is always true, no matter wat happens in the world ! so its an idenity !